# Loss Functions for Measuring the Accuracy of Nonnegative Cross-Sectional Predictions

### v2

@@ 6.5 Apportionment-Based Loss Functions
Spencer (1985) develops five apportionment-based loss functions. In addition to the Webster-Saint Lagüe Rule, these are (1) $p=1$, $q=0$, or more generally, $p>0,q=0$. (Hamilton’s method), (2) $p=2,q=-1$ with $P_{i}$ raised to the $q$ instead of $A_{i}$ (Hill’s method/Method of Equal Propertions), (3) $\max_{i}P_{i}/A_{i}$ (Jefferson/d’Hondt method) and (4) $\max_{i}A_{i}/P_{i}$ (Adam’s Method). Hamilton’s and Hill’s methods both violate Assumption 3. The Jefferson/d’Hondt method violates both Assumption 3 and the von Neumann-Morgenstern axioms. Adam’s Method additionally violates Assumptions 1 and 2. Spencer (1986, p. 398) advocates incorporating expectations into these methods to measure accuracy. However, since we have assumed the distributions generating the errors to be unknown, this is impossible.
@@ 6.6 Some Examples which Obey Assumptions 1-3 and Property 1
[…]
@@ 7 Loss Function-Based Bias Measures
where $S_{i}$ is interpreted as the signed loss function. Besides being a measure of bias, it can rank observations by how overprediction or underprediction can contributes to the loss of accuracy. Using it with external data may enable identification of factors affecting accuracy.

### v3

@@ 6.5 Apportionment-Based Loss Functions
Spencer (1985) develops five apportionment-based loss functions. In addition to the Webster-Saint Lagüe Rule, these are (1) $p=1,q=0$, or more generally, $p>0,q=0$ loss function. (Hamilton’s method), (2) $p=2,q=-1$ with $P_{i}$ raised to the $q$ instead of $A_{i}$ (Hill’s method/Method of Equal Propertions), (3) $\max_{i}P_{i}/A_{i}$ (Jefferson/d’Hondt method) and (4) $\max_{i}A_{i}/P_{i}$ (Adam’s Method). Hamilton’s and Hill’s methods both violate Assumption 3. The Jefferson/d’Hondt method violates both Assumption 3 and the von Neumann-Morgenstern axioms. Adam’s Method additionally violates Assumptions 1 and 2. Spencer (1986, p. 398) advocates incorporating expectations into these methods to measure accuracy. However, since we have assumed the distributions generating the errors to be unknown, this is impossible.
@@ 6.6 Some Examples which Obey Assumptions 1-3 and Property 1
Bryan (2000) proposes using $p=1$ and $q=\log(range)/25-1$ in the context of small-area population estimates evaluation, where $range$ is the range of the $A_{i}$. This obeys Assumption 3 and Property 1 for $range<\exp(25)\approx 72{,}000{,}000{,}000$. Apparently, $p =1$ and $q=-1/2$ was proposed for evaluating the U.S. Census Bureau’s Small Area Income and Poverty Estimates (William R. Bell, personal correspondence).
@@ 7 Loss Function-Based Bias Measures
where $S_{i}$ is interpreted as the signed loss function. Besides being a measure of bias, it can rank observations by how overprediction or underprediction can contributes to the loss of accuracy. Using it with external data may enable identification of factors affecting accuracy. [^35] Judson, Popoff and Batutis (1999) do this for the U.S. Census Bureau’s 1990 county population estimates using the Algebraic Percentage Error (ALPE) for their bias measure. ALPE is defined as $100*S_{i}(P,A)$ with $p=1$ and $q=-1$. Taking its mean also results in MALPE.